OpenStudy (anonymous):

The point (2, 3) is on the terminal side of angle θ, in standard position. What are the values of sine, cosine, and tangent of θ? Make sure to show all work.

OpenStudy (anonymous):

sdo you know what sine cosine and tangent mean

OpenStudy (anonymous):

sine = y cos=x \[sinx^2 + cosx^2 = r^2\]

OpenStudy (anonymous):

\[(2,3)\] \[2^2+ 3^2 = r^2\] \[\sqrt{13}= r\]

OpenStudy (anonymous):

\[sine \theta = \frac{ y }{ r }\] \[cos \theta = \frac{ x }{ r }\] \[tan \theta = \frac{ y}{ x }\] can you solve now?

OpenStudy (anonymous):

yes thank you :)

OpenStudy (anonymous):

\[\tan (\theta )=\frac{3}{2} \]\[\sin \left(\tan ^{-1}\left(\frac{3}{2}\right)\right)=\frac{3}{\sqrt{13}} \]\[\cos \left(\tan ^{-1}\left(\frac{3}{2}\right)\right)=\frac{2}{\sqrt{13}} \]